# Representational Analysis of Binding in Large Language Models

### v1

@@ A.1 Other Related Work
Knowledge Localization Many works aim to localize and edit factual relations (e.g., “capital of”) that LMs learn from pretraining and are stored into model weights Geva et al. (2020); Dai et al. (2021); Meng et al. (2022); Geva et al. (2023); Hernandez et al. (2023). Different from this line of research, this work studies in-context representations of relations and analyzes how they are stored in model activations.
Mechanistic Interpretability Notable progress has been made in uncovering circuits performing various tasks within LMs Elhage et al. (2021); Wang et al. (2022); Wu et al. (2024). Recently, Prakash et al. (2024) identify the circuit for entity tracking task. Feng and Steinhardt (2023) introduce a Binding ID Mechanism for explaining the binding problem, state that LMs use the abstract concept BI to internally mark entity-attribute pairs. However, they does not directly capture BI information from activations. Therefore, they have not answered how LMs store the BI information into entity activations, how to localize the BI information and whether the localized BI information causally affect the model binding behaviour.
@@ A.2 Partial least squares regression and PCA
Besides PCA, a commonly used unsupervised Dimension Reduction (DR) method, we also attempt Partial Least Squares regression (PLS) Wold et al. (2001), a supervised DR method. PLS extracts a set of ordered latent variables that maximizes the co-variability between the features (e.g., activations) and the scores to be predicted (e.g., BI). We perform PCA and PLS on a development set and compare their regression cures in Figure 11. We can observe that both the first PCA component and the first PLS direction contain almost all information about BI of target entity, because their regression score is close to one. The consistency indicates that PCA is an effective method to capture BI.
@@
A.3 Layer-wise Intervention
To determine how well BI subspace from different layers of a LM causally affects the model output, we perform layer-wise BI-PC based intervention mentioned in Section (§4.3) on our development set. In Figure 12, we can observe that BI subspace from middle layers (i.e., from layer7 to layer15, especially layer8) significantly affect the computation of binding, and interestingly, these layers also overlap with the ones that clearly encode BI information, as shown in Figure 2. Based on such analysis, we select the layer to perform activation patching.
Figure 12: Average Logit Difference (LD) and logit flip for layer-wise BI-PC based intervention on Llama2-7B, where x axis denotes the layer, the colored zone indicates the layers that are sensitive to the intervention, and the vertical line represents the most sensitive one (i.e., Layer 8), Y axis denotes the average LD and the proportion of inferred attributes (excluding the original one) in Figure 12(a) and Figure 12(b) respectively.
A.4 Layer-wise Embedding Visualization on Llama3-8B and Float-7B
@@
Figure 13: Layer-wise BI subspace visualization on Llama3-8B and Float-7B.
A.5 Layer-wise Embedding Visualization after Masking Binding Information
The counter argument is that the captured subspace only represents the positional information. To test the claim, we attempt to mask the context around the entities and attributes with random two-letter tokens (e.g., “td”) so as to ablate the binding information and keep the positional information. For instance, we convert “the document is in Box Q , the bus is in Box F, …” as “pl document td cy wa Q br fl bus ti eq fs F …” so that there is no relational information in the latter. Figure 14 compares the BI subspace distribution between with and without binding information. We can observe that ablation of binding information tangles the distribution so that there is no clear clustering for each BI (e.g., by comparing layer14). This indicates that our discovered subspace encodes binding information.
(a) w/ Binding Information
(b) w/o Binding Information
Figure 14: Layer-wise BI subspace visualization for w/ and w/o binding information on Llama2-7B.
A.6 Case Study on Llama2-7B
@@
Table 7: Attributes inferred by Llama2-7B as a result of directed activation patching along BI-PC in the BI subspace on the dataset of “r: sell”, where color denotes the BI.
@@
Table 8: Attributes inferred by Llama2-7B as a result of directed activation patching along BI-PC in the BI subspace on the dataset of “r: apply”, where color denotes the BI.
@@
Table 9: Attributes inferred by Llama2-7B as a result of directed activation patching along BI-PC in the BI subspace on the dataset of “r: move”, where color denotes the BI.
@@
Table 10: Attributes inferred by Llama2-7B as a result of directed activation patching along BI-PC in the BI subspace on the dataset of “r: bring”, where color denotes the BI.
@@
Table 11: Attributes inferred by Llama2-7B as a result of directed activation patching along BI-PC in the BI subspace on the dataset of “r: push”, where color denotes the BI.
A.7 Activation Steering on Llama2-7B
@@
Figure 15: Logit Difference (LD) for BI subspace based activation steering across datasets on Llama2-7B, where x axis represents the intervention of $\mathbf{s}_{0\rightarrow bi}$ on the activation of $e_{0}$. Here, $l=8$ and $\alpha=1.25$.
@@
Figure 16: Logit flip for BI subspace based activation steering across datasets on Llama2-7B, where x axis represents the intervention of $\mathbf{s}_{0\rightarrow bi}$ on the activation of $e_{0}$.
A.8 Activation Patching on Float-7B
@@
Figure 17: Logit Difference (LD) for BI-PC based intervention across datasets on Float-7B, where x axis denotes the number of intervention steps on $e_{0}$, y axis does the LD, BI_i represents each target attribute and the light yellow bottom line indicates the LD of original attribute (i.e., $a_{0}$). Here, $l=10$, $v=2.55$, and $\alpha=5.0$.
@@
Figure 18: Logit flip for BI-PC based intervention across datasets on Float-7B, where x axis denotes the number of intervention steps on $e_{0}$, y axis does the proportion of each inferred attribute in model output.
A.9 Activation Patching on Llama3-8B
@@
Figure 19: Logit Difference (LD) for BI-PC based intervention across datasets on Llama3-8B, where x axis denotes the number of intervention steps on $e_{0}$, y axis does the LD, BI_i represents each target attribute and the blue line indicates the LD of original attribute (i.e., $a_{0}$). Here, $l=10$, $v=0.65$, and $\alpha=2.0$.
@@
Figure 20: Logit flip for BI-PC based intervention across datasets on Llama3-8B, where x axis denotes the number of intervention steps on $e_{0}$, y axis does the proportion of each inferred attribute in model output.
A.10 Activation Patching on the New Dataset with Interjections

### v3

@@ A.1 Other Related Work
Knowledge Localization Many works aim to localize and edit factual relations (e.g., “capital of”) that LMs learn from pretraining and are stored into model weights Geva et al. (2020); Dai et al. (2021); Meng et al. (2022); Geva et al. (2023); Hernandez et al. (2023). Different from this line of research, this work studies in-context representations of relations and analyzes how they are represented in model activations.
Mechanistic Interpretability Notable progress has been made in uncovering circuits performing various tasks within LMs Elhage et al. (2021); Wang et al. (2022); Wu et al. (2024). Recently, Prakash et al. (2024) identify the circuit for entity tracking task. Feng and Steinhardt (2023) introduce a Binding ID Mechanism for explaining the binding problem, state that LMs use the abstract concept BI to internally mark entity-attribute pairs. However, they have not captured the OI information from entity activations that directly determines the binding behaviour.
@@ A.2 Partial least squares regression and PCA
Besides PCA, a commonly used unsupervised Dimension Reduction (DR) method, we also attempt Partial Least Squares regression (PLS) Wold et al. (2001), a supervised DR method. PLS extracts a set of ordered latent variables that maximizes the co-variability between the features (e.g., activations) and the scores to be predicted (e.g., OI). We perform PCA and PLS on a development set and compare their regression cures in Figure 11. We can observe that both the first PCA component and the first PLS direction contain almost all information about OI of target entity, because their regression score is close to one. The consistency indicates that PCA is an effective method to capture OI subspace.
@@
A.3 ICA for OI Subspace
Independent Component Analysis (ICA) is a statistical method to reveal hidden subcomponent from multivariate signal. Similar to PCA, ICA is utilized as a dimension reduction method. We apply ICA to analyze the activations of entities and visualize the subspace in Figure 12. We can observe that the distribution is generally similar to the one of PCA. In addition, we also conduct causal intervention (i.e., activation steering) based on the subspace and present the results in Figure 13. We can observe that the results are also similar to those from PCA based intervention. This indicates that besides PCA, ICA can also be used as a computational method to capture OI subspace.
Figure 12: Subspace visualization from ICA on Llama2-7B, where “BI” primarily denotes OI.
Figure 13: Logit Difference and Logit Flip for ICA based activation steering on the entity tracking dataset (i.e., r: is_in) on Llama2-7B.
A.4 Layer-wise Intervention
To localize the layer that contributes to binding behaviour, we perform layer-wise OI-PC based intervention mentioned in Section (§4.3) on our development set. In Figure 14, we can observe that OI subspace from middle layers (i.e., from layer7 to layer15, especially layer8) significantly affect the computation of binding, and interestingly, these layers also overlap with the ones that clearly encode OI information, as shown in Figure 2. Based on the analysis, we select the layer to perform activation patching.
Figure 14: Average Logit Difference (LD) and logit flip for layer-wise OI-PC based intervention on Llama2-7B, where x axis denotes the layer, the colored zone indicates the layers that are sensitive to the intervention, and the vertical line represents the most effective layer (i.e., Layer 8), Y axis in Figure 14(a) and Figure 14(b) denotes the average LD and the proportion of inferred attributes (excluding the original one) respectively.
A.5 Layer-wise Embedding Visualization on Llama3-8B and Float-7B
@@
Figure 15: Layer-wise OI subspace visualization on Llama3-8B and Float-7B.
A.6 OI Subspace on other LM Families
(a) Qwen1.5-7B
(b) Pythia-6.9B
Figure 16: Layer-wise OI subspace visualization on Qwen1.5-7B and Pythia-6.9B.
(a) LD for Qwen1.5-7B.
(b) LF for Qwen1.5-7B.
(c) LD for Pythia-6.9B.
(d) LF for Pythia-6.9B.
Figure 17: Logit Difference (LD) and Logit Flip (LF) for activation patching on the entity tracking dataset (i.e., r: is_in).
A.7 Case Study on Llama2-7B
@@
Table 7: Attributes inferred by Llama2-7B as a result of directed activation patching along OI-PC on the dataset of “r: sell”, where color denotes the BI.
@@
Table 8: Attributes inferred by Llama2-7B as a result of directed activation patching along OI-PC on the dataset of “r: apply”, where color denotes the BI.
@@
Table 9: Attributes inferred by Llama2-7B as a result of directed activation patching along OI-PC on the dataset of “r: move”, where color denotes the BI.
@@
Table 10: Attributes inferred by Llama2-7B as a result of directed activation patching along OI-PC on the dataset of “r: bring”, where color denotes the BI.
@@
Table 11: Attributes inferred by Llama2-7B as a result of directed activation patching along OI-PC on the dataset of “r: push”, where color denotes the BI.
A.8 Activation Steering on Llama2-7B
@@
Figure 18: Logit Difference (LD) for OI subspace based activation steering across datasets on Llama2-7B, where x axis represents the intervention of $\mathbf{s}_{0\rightarrow bi}$ on the activation of $e_{0}$. Here, $l=8$ and $\alpha=1.25$.
@@
Figure 19: Logit flip for OI subspace based activation steering across datasets on Llama2-7B, where x axis represents the intervention of $\mathbf{s}_{0\rightarrow bi}$ on the activation of $e_{0}$.
A.9 Results on Fine-Tuned LM
Kramár et al. (2024); Kim et al. (2024) claim that the code fine-tuned LM, such as Float-7B Prakash et al. (2024) outperforms the pretrained LM on the entity tracking task Kim and Schuster (2023); Prakash et al. (2024). Since the code fine-tuned LM performs well on the entity tracking task that requires the OI subspace based computation, we hypothesize that OI subspace also exists in the code fine-tuned LM and the intervention along OI-PC will causally affect the model output. To prove the hypothesis, we conduct the intervention on Float-7B and show results in Figure 20 and Figure 21. We found that the OI subspace based intervention on Float-7B achieves the similar results as on Llama2-7B, indicating that the OI subspace not only exists in the pretrained LM but also in the fine-tuned one. In addition, adding the same step value (i.e., $v$) on Float-7B will achieve higher LD value than Llama2-7B, indicating that the code fine-tuned LM is more sensitive to the OI subspace based intervention. For instance, the maximum LD of $a_{4}$ in the former is around $10$, and it is $2$ times larger than the one in the latter, which is around $5$. This might partially explains why the code fine-tuned LM performs better than the original one, because code fine-tuning might enhance the function of OI subspace so that it is more sensitive on the intervention and more effective on the in-context entity tracking task.
@@ A.9 Results on Fine-Tuned LM
Figure 20: Logit Difference (LD) for OI-PC based intervention across datasets on Float-7B, where x axis denotes the number of intervention steps on $e_{0}$, y axis does the LD, BI_i represents each target attribute and the light yellow bottom line indicates the LD of original attribute (i.e., $a_{0}$). Here, $l=10$, $v=2.55$, and $\alpha=5.0$.
@@
Figure 21: Logit flip for OI-PC based intervention across datasets on Float-7B, where x axis denotes the number of intervention steps on $e_{0}$, y axis does the proportion of each inferred attribute in model output.
A.10 Activation Patching on Llama3-8B
@@
Figure 22: Logit Difference (LD) for OI-PC based intervention across datasets on Llama3-8B, where x axis denotes the number of intervention steps on $e_{0}$, y axis does the LD, BI_i represents each target attribute and the blue line indicates the LD of original attribute (i.e., $a_{0}$). Here, $l=10$, $v=0.65$, and $\alpha=2.0$.
@@
Figure 23: Logit flip for OI-PC based intervention across datasets on Llama3-8B, where x axis denotes the number of intervention steps on $e_{0}$, y axis does the proportion of each inferred attribute in model output.
A.11 New Dataset with Interjections
The dataset is created by inserting a sequence of interjections after the first attribute entity pair, as illustrated in Table 12. Since there is no BI information in the interjection (e.g., $j^{p3}$), adding it only changes the PI of its following entities and attributes. We set the number of interjections as that the PI of last interjection token is larger than the last PI of its original input (e.g., $pi>p5$ in Table 12).
Based on this dataset, we conduct the same intervention on its OI subspace, as mentioned in Section (§4.3). The counter argument is that the subspace only captures Position Information (PI), and the intervening step only changes the PI information. Specifically, adding one unit of $v$ on $\overrightarrow{e_{0}^{p1}}$ might convert the PI of target entity from $p1$ to $p3$, and $p3$ is the PI of $e_{1}$, and its attribute is $a_{1}$, as shown in Table 12, and thus the LM swaps the answer from $a_{0}$ to $a_{1}$. If it is true, then the same intervention will not change the answer on the new dataset, because following the counter argument, after adding one unit of $v$ on $\overrightarrow{e_{0}^{p1}}$, the PI of target entity becomes $p3$, and $p3$ is the PI of $j^{p3}$ (i.e., an interjection token). The LM thus would not select $a_{1}$ as its answer. However, the results on Figure 24 and Figure 25 show that the subspace intervention on the new dataset achieves similar results as the original one, as shown in Figure 3 and Figure 4, proving the counter argument wrong and indicating the independence between OI subspace and PI.
Input (original)
$a_{0}^{p0}\,r\,e_{0}^{p1},\,a_{1}^{p2}\,r\,e_{1}^{p3},\,a_{2}^{p4}\,r\,e_{2}^{p5}.\quad e_{0}^{p1}\ r^{-1}\ ?$
Input (with interjection)